OpenStudy (anonymous):

Find an equation of a plane. The plane passes through the point(-1,2,1) and contains the line of intersection of the planes x + y - z = 2 and 2x - y +3z = 1 help please

OpenStudy (anonymous):

Well, first you need to find the line where the two planes meet, can you do this?

OpenStudy (anonymous):

i am not sure

OpenStudy (anonymous):

So far I took the normals of the 2 planes and did the cross product where I got <2, -5, -3> this is the direction vector. I found a point by making z = 0 for both equations of the plane. My newly found point is (1,1,0) I plug this new point 2(x-1) -5(y-1) -3(z-0) =0 therefore 2x -5y -3z = -3 and this is not what they get

OpenStudy (amistre64):

if you cross the normals, i believe you get the vector for the line

OpenStudy (amistre64):

using the normal vector <a,b,c> and a point common to both planes (its on the line of intersection) we can form the line equation as:\[x=x_o+at\]\[y=y_o+bt\]\[z=z_o+ct\]

OpenStudy (amistre64):

you tried z=0 to find a common point x + y = 2 2x - y = 1 ---------- 3x = 3 ; x=1 when x=1, y=1 therefore the point (1,1,0) is common to both planes

OpenStudy (amistre64):

lets see how your cross worked out x 1 2 x: 3-1 = 2 y 1 -1 -y: 3+2 = -5 z -1 3 z:-1-2 = -3 that was good

OpenStudy (amistre64):

sooo the equation for the line is \[x=1+2t\]\[y=1-5t\]\[z=-1-3t\]

OpenStudy (amistre64):

ugh, i used a vector as a point lol ... lets try that with the 1,1,0

OpenStudy (amistre64):

\[x=1+2t\]\[y=1-5t\]\[x=0-3t\] thats better

OpenStudy (amistre64):

...and i see that im reading the final outcome wrong; its not asking for the line equation is it; but the plane that contains the vector and a given point.

OpenStudy (anonymous):

Thats is right is has me going in circles

OpenStudy (amistre64):

well, we only need 2 vectors to define a normal for the plane that contains them right?

OpenStudy (amistre64):

one vector is <2,-5,-3> the other is the vector from (1,1,0) to (-1,2,1)

OpenStudy (anonymous):

so with our new found point where extract the other vector = <-2,1,1> then do cross product with <2,-5,-3>

OpenStudy (amistre64):

(1,1,0) -(-1,2,1) -------- <2,-1,-1> ; or as you did <-2,1,1> yes cross them and attach that normal to either point for the setup of the plane

OpenStudy (anonymous):

The X product i get is <2,-4,8> then drop in any point say (-1,2,1) they get as final answer x -2y +4z = -1 which is different to where we are heading

OpenStudy (amistre64):

<2,-4,8>; or simplified as: <1,-2, 4> \[1(x-x_o)-2(y-y_o)+4(z-z_o)=0\] \[(x-1)-2(y-1)+4(z-0)=0\] \[x-1-2y+2+4z=0\] \[plane:~x-2y+4z=-1\] where are we heading?

OpenStudy (anonymous):

That is brilliant, so you factored out 2. how do I give you a medal on this site?

OpenStudy (amistre64):

to the right of my name there should be a "best response" option that you can click. if it is not there, you might have to refresh the screen.

OpenStudy (anonymous):

Thanks very much for your help

OpenStudy (amistre64):

youre welcome, and good luck ;)

OpenStudy (anonymous):

What textbook are you using to get this stuff? I am using James Stewart Calculus